OpenStudy (anonymous):

Put this in partial fractions: \[1 \over y(4-y)\]

OpenStudy (anonymous):

@NotSObright

OpenStudy (anonymous):

\[\frac{1}{y(4-y)}=\frac{A}{y}+\frac{B}{4-y}\]

OpenStudy (anonymous):

Combine fractions on right hand side so that you will have same denominator => numerators must be same

OpenStudy (anonymous):

\[\frac{1}{y(4-y)}=\frac{A(4-y)+By}{y(4-y)}\]

OpenStudy (anonymous):

\[1=A(4-y)+By\]

OpenStudy (anonymous):

\[1=4A-Ay+By\] \[1=4A+y(-A+B)\]

OpenStudy (anonymous):

That's where I got... don't know how to carry on

OpenStudy (anonymous):

\[4A=1 , -A+B=0\]

OpenStudy (anonymous):

Constants on both side set them equal set the coefficients of y on both sides equal and so on...

OpenStudy (anonymous):

Multiply both sides by y and let y=0 \[ \frac{1}{y(4-y)}=\frac{A}{y}+\frac{B}{4-y}\\ A= \frac 1 4 \] Multiply both sides by4- y and let y=4 \[ \frac{1}{y(4-y)}=\frac{A}{y}+\frac{B}{4-y}\\ B=- \frac 1 4 \]

OpenStudy (anonymous):

This is called the cover-up method.

OpenStudy (anonymous):

Why aren't they both 1/4? With Zombie's method, they'd both equal 1/4

OpenStudy (anonymous):

Sorry \[B= \frac 1 4 \]

OpenStudy (anonymous):

OK Thanks :)